OpenStudy (anonymous):

from 1 to 9 1/(x-9)^1/3 Determine whether the integral is convergent or divergent.

OpenStudy (anonymous):

Is it divergent? Because when x is 9 it isn't a real number. Right?..

OpenStudy (sidsiddhartha):

have u integrated the function?

OpenStudy (sidsiddhartha):

\[\large \int\limits_{}^{}\frac{ 1 }{ \sqrt[3]{x-9} }=\frac{ 3 }{ 2 }*(x-9)^{2/3}\] now put the limits

OpenStudy (anonymous):

I think so, I got 3(x-9)^(3/2)/2

OpenStudy (anonymous):

it converges at 6??

OpenStudy (anonymous):

I need a value it converges to

OpenStudy (anonymous):

I used u = x - 9.. it is improper at u = 0, so I plugged values back in for 'u' to find the new bounds. For 9 u is 0... for 1 u is -8. Is this right? Please help

OpenStudy (sidsiddhartha):

if the integral have a finite value between the limits then it will converge

OpenStudy (anonymous):

So is this right? integrate from -8 to 0. I am so confused.

OpenStudy (anonymous):

medal for whoever helps me solve this D:

OpenStudy (sidsiddhartha):

yes limits will be -8 to 0

OpenStudy (anonymous):

okay. I got 9.917234

OpenStudy (anonymous):

is this what you got too?

OpenStudy (anonymous):

I just want to learn how to solve this so I don't spend so much time do it wrong

OpenStudy (sidsiddhartha):

so it produces a finite value then it should converge

OpenStudy (anonymous):

yes, but I need the exact value. Is this right?

OpenStudy (anonymous):

my question is, what IS the value it converges at

OpenStudy (anonymous):

\[\int_1^9\frac{dx}{(x-9)^{1/3}}\] Substitute \(u=x-9\), so \(du=dx\) and you have \[\int_{-8}^0\frac{du}{u^{1/3}}\] The integral is improper, so treat as a limit: \[\lim_{c\to0^+}\int_{-8}^c\frac{du}{u^{1/3}}\] Power rule: \[\lim_{c\to0^+}\left[\frac{3u^{2/3}}{2}\right]_{-8}^c\] Fundamental theorem: \[\frac{3}{2}\lim_{c\to0^+}\left[c^{2/3}-(-8)^{2/3}\right]=\frac{3}{2}\left[0-(-8)^{2/3}\right]=\frac{3}{2}(-4)=-6\]

OpenStudy (anonymous):

Should be POSITIVE 6, right? Double negatives. Just making sure.